A man made satellite of mass 400 kg orbits the Earth with a radius of 8.2 x 10^6 m. Linear speed of the satellite is 6.96 x 10^3 m/s. What is the centripetal force acting on the satellite?



1
Expert's answer
2021-05-27T08:31:16-0400
"F=\frac{mv^2}{r}\\F=\frac{(400)(6960)^2}{8.2\cdot10^6}=2363\ N"